Herschel Status Report - May 2010
The first major scientific symposium presenting Herschel results took place from 4 to 7 May 2010 in ESTEC, Noordwijk, The Netherlands. The presentations and the available posters from this 'Herschel First Results Symposium' (ESLAB 2010) are posted on the Herschel Science Centre website (see right-hand menu).
The spacecraft is operating nominally.
Star tracker warm pixels
There is now a thorough analysis by the Herschel Science Centre of the effects of having lowered the star tracker CCD temperature to -10 °C. The lower CCD temperature was introduced to limit the occurrences of transient "warm" pixels (see status report from April 2010). These pixels carry additional charge but appear and disappear on a timescale of hours to many days. They can result in inhomogeneous scan speeds ("speed bumps") along some scan lines when a guide star crosses them during a mapping observation (see also previous status reports, back to October 2009).
Since the lowering of the CCD temperature, only a single "speed bump" has occurred and it was also of a smaller magnitude. In the past it would have been expected to see significantly more than 100 "speed bumps" over a comparable time period. Thus, the situation is now fully stable from an operational point of view.
Two further measures to stay ahead of the problem - should it be one of gradual degradation over the lifetime of the star tracker CCD under in-orbit conditions - are currently being prepared:
- A further reduction in the CCD temperature. The Peltier cooler for the star tracker CCD is believed to easily reach temperatures of -25 °C and below, but a reduction to below the current CCD temperature (-10 °C) requires a change in the on-board software, which needs to be carefully considered and tested.
- In-flight updates of the star tracker "bad pixel" table based on the findings of the full CCD dumps that are performed every other week.
All three instruments, HIFI, PACS and SPIRE, continue to operate nominally. They have all entered the routine phase with only one more observing mode (HIFI on-the-fly mapping) left to be released.
Since mid-January 2010, when HIFI was switched back on using the redundant branch of its signal electronics, 9 bit-flips have been observed on 7 different occasions. Although they are still in the process of being refined, the operational procedures to recover from such events have proven to be robust and reliable in preventing any damage to this instrument.
During the reporting period, mission operations have been conducted with the support of the ESA New Norcia ground station. Observational data stored on-board Herschel was received on ground during communication passes lasting approximately three hours.
The ground segment is operating nominally. Data products are generated routinely and ingested into the Herschel Science Archive (HSA).
- 20 May 2010: Issuing of first in-flight Announcement of Opportunity (AO) for Open Time (OT1) observations
- 22 July 2010: Closing date of first in-flight AO for OT1 observations
- 1 November 2010: Announcement of approved OT1 proposals